OpenStudy (anonymous):

What are the coordinates of the vertex of the function below? Write your answer in the form (x,y). y-11= -(x-5)^2

OpenStudy (anonymous):

@LaceyLeanne Do you know how to do this? :) I know I would use the equation y=a(x-h)^2+k but im not too sure how to replace the letters with the given numbers :O

OpenStudy (anonymous):

Okay so, What do you think you should do first?

OpenStudy (anonymous):

here is an example that i found A) THE SIMPLETST way to solve for the x-coordinates of your function is to simply set each factor equals to 0 and solve, like this: y=(x-5)(x+3) x - 5 = 0 x = 5 x + 3 = 0 x = -3 Thus, the x-coordinates of the vertex of your function are: (5, 0) (-3, 0) B) Here are some very useful general characteristics of a quadratic function and its graph:: 1. The graph of a quadratic function f(x) = ax+ bx + c is called a parabola. 2. If a > 0, the parabola opens upward; if a < 0, the parabola opens downward. 3. As | a| increases, the parabola becomes narrower; as | a| decreases, the parabola becomes wider. 4. The lowest point of a parabola (when a > 0) or the highest point (when a < 0) is called the vertex. 5. To determine the vertex of the graph of a quadratic function, f(x) = ax+ bx + c, we can either: ...a) use the method of completing the square to rewrite the function in the form f(x) = a(x – h) + k. The vertex is (h, k). , or ...b) use the formula x = -b / 2a to find the x-coordinate of the vertex Source(s):

OpenStudy (anonymous):

welll i think i would foil the x-5^2 right?

OpenStudy (anonymous):

Lacey is a math wiz

OpenStudy (anonymous):

i just give people examples lol

OpenStudy (anonymous):

hahah well thank you for the example :) but im still lost ._.

OpenStudy (anonymous):

lol soz

OpenStudy (anonymous):

Okay IceIce Thanks but hit the road, your confusing Ali!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you for helping out ice :)

OpenStudy (anonymous):

np but let the math wiz take it over @LaceyLeanne

OpenStudy (anonymous):

So are you trying to find the vertex?

OpenStudy (anonymous):

yah, it says to find the vertex and to put it in coordinates

OpenStudy (anonymous):

how would i do this @LaceyLeanne ?

OpenStudy (anonymous):

I am so sorry! The internet is messing up.

OpenStudy (anonymous):

its okay! do you know how i would solve this type of problem? :o

OpenStudy (anonymous):

Okay hold on Ali....My gawd I feel so bad for not helping you right away! Forgive me :)

OpenStudy (anonymous):

of course, i completely understand! :)

OpenStudy (anonymous):

Okay it is going to be negative because the slop is going down from the zero...Do you understand?

OpenStudy (anonymous):

slope*

OpenStudy (anonymous):

It appears to be -14

OpenStudy (anonymous):

yes, because the slope is going down instead of up

OpenStudy (anonymous):

That would make it a negative...Hold on let me graph it....

OpenStudy (anonymous):

Yes (0,-14)

OpenStudy (anonymous):

so those coordinates would be the answer?

OpenStudy (anonymous):

oh i see! that makes sense!

OpenStudy (anonymous):

Yeah, Thanks for being patient with this slow gurl :P lol

OpenStudy (anonymous):

anytime! thank you for helping me! i really appreciate more than you understand! :)

OpenStudy (anonymous):

No Problem, I am always a PM away.